# Consistent and Invariant Generalization Learning for Short-video Misinformation Detection

### v2

@@ Abstract.
Short-video misinformation detection has attracted wide attention in the multi-modal domain, aiming to accurately identify the misinformation in the video format accompanied by the corresponding audio. Despite significant advancements, current models in this field, trained on particular domains (source domains), often exhibit unsatisfactory performance on unseen domains (target domains) due to domain gaps. To effectively realize such domain generalization on the short-video misinformation detection task, we propose deep insights into the characteristics of different domains: (1) The detection on various domains may mainly rely on different modalities (i.e., mainly focusing on videos or audios). To enhance domain generalization, it is crucial to achieve optimal model performance on all modalities simultaneously. (2) For some domains focusing on cross-modal joint fraud, a comprehensive analysis relying on cross-modal fusion is necessary. However, domain biases located in each modality (especially in each frame of videos) will be accumulated in this fusion process, which may seriously damage the final identification of misinformation. To address these issues, we propose a new DOmain generalization model via ConsisTency and invariance learning for shORt-video misinformation detection (named DOCTOR), which contains two characteristic modules: (1) We involve the cross-modal feature interpolation to map multiple modalities into a shared space and the interpolation distillation to synchronize multi-modal learning; (2) We design the diffusion model to add noise to retain core features of multi modal and enhance domain invariant features through cross-modal guided denoising. Extensive experiments demonstrate the effectiveness of our proposed DOCTOR model.  [^1] Our code is publicly available at https://github.com/ghh1125/DOCTOR
@@ 1. INTRODUCTION
Figure 2. The problem of Domain-bias Enhancement in Cross-modal Fusion. The red line is a simple case of bias accumulated.
To tackle the aforementioned issues, we introduce a new DOmain generalization model via ConsisTency and invariance learning for shORt-video misinformation detection (named DOCTOR). Specifically, there are two specific designs: (1) Cross-modal Interpolation Distillation. We project video and audio modality features into a shared representation space, and mix these representations with a random mixing ratio to construct interpolations. Based on the feature distribution of these interpolation samples as teacher signals, feature distillation drives the learning optimization process of the student model, making the representation learning across modalities more consistent and synchronized. (2) Multi-modal Invariance Fusion. Incorporating subtitles extracted from each video frame as a guide, we devise a novel mechanism to facilitate the extraction of core video features from both spatial and temporal viewpoints. With extracted multi-modal features, we add noise to them step by step based on the diffusion model to filter out redundant details, thereby underlining the core features of multi-modal. Subsequently, by gradually mitigating the noise influence, we further outline the fundamental information of multi-modal features through cross-modal mutual guidance, boosting domain invariant features.
The main contributions of our work are as follows:
@@
Figure 3. The architecture of the DOCTOR model. (a) Cross-modal Interpolation Distillation: Video and audio modality features are projected into a shared representation space and mixed with a random ratio to create interpolation samples. These samples are then used as teacher signals to guide the student model in learning consistent representations across modalities. (b) Cross-modal Invariance Fusion: Multi-modal features are first fused, and then progressively perturbed through a diffusion model to mask domain biases. The noise is gradually removed to enhance the core features of multi-modal, with the goal of obtaining domain-invariant features.
@@ 2.1. Multi-modal Misinformation Detection
With the rise of online short-video platforms, the spread of misinformation has expanded to video-audio. Some short-video misinformation detection methods have been proposed through recent research for these situations. OpEvFake (Zong et al., 2024) employs a large language model to mine implicit viewpoints in short videos and combines model information to enhance interaction for judgment. SV-FEND (Qi et al., 2023) exploits cross-modal correlations to select the most informative features and uses social context information for detection. FakingRecipe (Bu et al., 2024) detects misinformation by simulating the creation process of audio and video. TikTec (Shang et al., 2021) employs speech text-guided visual object features and MFCC-guided speech textual features. However, these methods are typically designed for general domains and often underperform in domain generalization scenarios.
@@ 3. PROPOSED METHOD: DOCTOR
In this section, we propose an innovative DOmain generalization model via ConsisTency and invariance learning for shORt-video misinformation detection (named DOCTOR). We will present each module and its training strategy in detail as follows.
@@ 3.1.2. Overview of DOCTOR
As illustrated in Figure 3, which demonstrates the domain generalization for the short-video misinformation detection task, the proposed model, DOCTOR, is designed to address the generalization gap between source and target domains.
@@ 3.2. Cross-modal Interpolation Distillation
If the loss of mixed representation $\hat{X}_{a,v}$ can be optimized to lower values, we can achieve the best state in all modalities at the same time. However, the amount of computation in multi-modal learning is often huge, and it is often difficult to achieve the optimal state, which is impractical in practical applications. Thus, we adopt the cross-modal distillation technique to distill the knowledge in the mixed representation into each modality and then optimize the learned representation to achieve the best state in all modalities. Specifically, as follows, for each modal $k\in{(a:Audio,v:Video)}$, we initialize the student model $\gamma_{t}^{k}$ and utilize Simple Moving Average (SMA) to create the teacher model $\hat{\gamma_{t}^{k}}$, which is used to generate more stable and generalizable representations. The $\hat{\gamma}_{t}^{k}$ represents as:
@@
where $\gamma_{t}^{k}$ means that iteration $t$ is the student mode state of modal $k$, and $t_{0}$ is the number of iterations at which to start the SMA. Therefore, the representation produced by the teacher model $\hat{\gamma}_{t}^{k}$ is denoted as $\dot{X}_{k}$, which serves as a stable reference for the modality $k$. Based on this, we obtain a new cross-modal representation mixup $\dot{X}_{a,v}$ as follows:
@@
where $B$ is the sample set in the training batch, $\tau$ is the temperature coefficient, $\exp(\cdot/\tau)$ is a similarity scaling function used to adjust the discrimination between positive and negative samples, and $\hat{X}_{a}^{i},\hat{X}_{v}^{i}$ are the audio and video modality embedding vectors of sample $i$.
@@ 3.3. Cross-modal Invariance Fusion
Toward this issue, we propose an innovative method: Cross-modal Invariance Fusion. It designs a novel mechanism for performing core extraction for video features from spatial and temporal perspectives. With extracted multi-modal features, we add noise to them step by step based on the diffusion model to filter out redundant details, thereby underlining the core features of multi-modal. Subsequently, by gradually mitigating the noise influence, we further outline the core information of multi-modal features through cross-modal mutual guidance, enhancing domain invariant features. We aim to mitigate the impact of domain bias and retain domain invariance to enhance the domain generalization ability for misinformation detection. In detail, it includes the following two steps:
3.3.1. Multimodal Feature Fusion
@@
(9) $X_{time}=\text{MEAN}(\text{Self-Att}([Seg^{1},Seg^{2},...,Seg^{i}]).$
3.3.2. Diffusion Model Filtering Bias
@@
In the forward diffusion process, we denote the initial distribution of each fusion feature as $X_{emo}^{0}$ (emotion feature), $X_{sem}^{0}$ (semantic feature), $X_{spa}^{0}$ (spatial feature), $X_{time}^{0}$ (temporal feature). For each feature, we gradually incorporate noise sampled from a Gaussian distribution into the true distribution. The steps are as follows:
@@
where the equation follows the Markov chain principle and represents the distribution of $X_{F}^{k}$ under $X_{F}^{k-1}$, $\beta_{F}^{k}$ is a weight and indicates the proportion of noise added at the $k$-th step. $I$ represents the identity matrix, which is used to control the covariance structure of the noise. $F$ is a different feature of emo feature, sem feature, spa feature, and time feature. $X_{F}^{k}$ can be sampled through the following process:
@@
where $\epsilon$ is the noise sampled from $\mathcal{N}(0,I)$. The overall noise injection process is as follows, effectively filtering out redundant biases:
@@
where $\mu_{\theta}$ is the predicted mean (indicating the direction of denoising), and $\Sigma_{\theta}(k)$ is the covariance of the noise. This process gradually generates $X_{F}^{k-1}$, $X_{F}^{k-2}$, … , $X_{F}^{0}$ from $X_{F}^{k}$ that is, gradual denoising.
4. EXPERIMENTS
4.1. Experimental Setups
4.1.1. Datasets
4.1.2. Baselines
We selected some popular misinformation short-video detection methods, such as FakingRecipe (Bu et al., 2024), OpEvFake (Zong et al., 2024), TikTec (Shang et al., 2021), HCFC-Hou (Hou et al., 2019), HCFC-Medina (Serrano et al., 2020), FANVM (Choi and Ko, 2021), SV-FEND (Qi et al., 2023). We also compared with the novel general domain generalization method CMRF (Fan et al., 2024) and the text-image misinformation domain generalization methods MMDFND (Tong et al., 2024) and MDFEND (Nan et al., 2021). We also select GPT-4 (Achiam et al., 2023) and GPT-4V (Yang et al., 2023b) for comparison.
4.1.3. Implementations
All of our experiments were conducted on an NVIDIA RTX A6000. More information see code link.
4.2. Experimental Results
4.2.1. Domain generalization experiments
Table 2. Comparison of general domain misinformation detection performance of different methods on FakeSV and FakeTT datasets.
@@ 4.2.1. Domain generalization experiments
Table 3. Learning rate for DOCTOR model (Accuracy).
Dataset 1e-5 5e-5 1e-4 5e-4 1e-3 5e-3
FakeSV 82.32 84.91 83.45 80.78 79.53 78.44
FakeTT 72.24 74.31 76.86 78.57 78.92 77.39
@@ 4.2.4. Ablation experiments
Figure 4. T-SNE visualizations display the different modal, cross-modal fusion and DOCTOR model. The contents of Figure are as follows: (a) Single frame of video, (b) Video (all frames), (c) Audio, (d) Video-Audio fusion, (e) DOCTOR model (Ours).
4.2.5. Visualization results
Due to the domain bias present in videos from different domains, simple fusion techniques may inadvertently amplify this bias. To address this issue, DOCTOR introduces Cross-modal Invariance Fusion to mitigate the impact of domain bias while preserving domain-invariant features. This approach enhances the model’s domain generalization capability for misinformation detection. Additionally, we utilized t-SNE (Van der Maaten and Hinton, 2008) to visualize the features of different modalities and their fusion. As shown in Figure 4. We can find that DOCTOR can effectively reduce domain bias through Cross modal Invariance Fusion.
@@ 5. CONCLUSION
In this paper, we explore the shortcomings of the domain generalization of short-video misinformation detection, which are mainly divided into two aspects: (1) Varying Modal-dependence in Different Domains and (2) Domain-bias Enhancement in Cross-modal Fusion. To bridge these gaps, we propose an innovative DOmain generalization model via ConsisTency and invariance learning for shORt-video misinformation detection (named DOCTOR). It incorporates two main modules: (a) Cross-modal Interpolation Distillation and (b) Cross-modal Invariance Fusion. We validate the domain generalization effectiveness of misinformation detection of DOCTOR through comprehensive comparative experiments. In addition, in-depth analysis, including ablation studies, demonstrates the design rationality of each module in the DOCTOR model.
Acknowledgment

### v3

@@ Abstract.
Short-video misinformation detection has attracted wide attention in the multi-modal domain, aiming to accurately identify the misinformation in the video format accompanied by the corresponding audio. Despite significant advancements, current models in this field, trained on particular domains (source domains), often exhibit unsatisfactory performance on unseen domains (target domains) due to domain gaps. To effectively realize such domain generalization on the short-video misinformation detection task, we propose deep insights into the characteristics of different domains: (1) The detection on various domains may mainly rely on different modalities (i.e., mainly focusing on videos or audios). To enhance domain generalization, it is crucial to achieve optimal model performance on all modalities simultaneously. (2) For some domains focusing on cross-modal joint fraud, a comprehensive analysis relying on cross-modal fusion is necessary. However, domain biases located in each modality (especially in each frame of videos) will be accumulated in this fusion process, which may seriously damage the final identification of misinformation. To address these issues, we propose a new DOmain generalization model via ConsisTency and invariance learning for shORt-video misinformation detection (named DOCTOR), which contains two characteristic modules: (1) We involve the cross-modal feature interpolation to map multiple modalities into a shared space and the interpolation distillation to synchronize multi-modal learning; (2) We design the diffusion model to add noise to retain core features of multi modal and enhance domain invariant features through cross-modal guided denoising. Extensive experiments demonstrate the effectiveness of our proposed DOCTOR model. Our code is publicly available at https://github.com/ghh1125/DOCTOR.
@@ 1. Introduction
Figure 2. The problem of domain-bias Enhancement in Cross-modal Fusion. The red line is a case of bias accumulated.
@@
To tackle the aforementioned issues, we introduce a new DOmain generalization model via ConsisTency and invariance learning for shORt-video misinformation detection (named DOCTOR). Specifically, there are two specific designs: (1) Cross-modal Interpolation Distillation. We project video and audio modality features into a shared representation space, and mix these representations with a random mixing ratio to construct interpolations. Based on the feature distribution of these interpolation samples as teacher signals, feature distillation drives the learning optimization process of the student model, making the representation learning across modalities more consistent and synchronized. (2) Multi-modal Invariance Fusion. Incorporating subtitles extracted from each video frame as a guide, we devise a novel mechanism to facilitate the extraction of core video features from both spatial and temporal viewpoints. With extracted multi-modal features, we add noise to them step by step based on the diffusion model to filter out redundant details, thereby underlining the core features of multi-modal. Subsequently, by gradually mitigating the noise influence, we further outline the fundamental information of multi-modal features through cross-modal mutual guidance, boosting domain invariant features. The main contributions of our work are as follows:
@@
Figure 3. The framework of the DOCTOR. (a) Cross-modal Interpolation Distillation: Video and audio features are projected into a shared representation space and mixed with a random ratio to create interpolation samples. These samples are then used as teacher signals to guide the student model in learning consistent representations across modalities. (b) Cross-modal Invariance Fusion: Multi-modal features are first fused, then progressively perturbed through a diffusion model to mask domain biases. The noise is gradually removed to enhance core multi-modal features, aiming to obtain domain-invariant features.
@@ 2.1. Multi-modal Misinformation Detection
With the rise of online short-video platforms, the spread of misinformation has expanded to video-audio. Some short-video misinformation detection methods have been proposed through recent research. OpEvFake (Zong et al., 2024) employs a large language model to mine implicit viewpoints in short videos and combines model information to enhance interaction for judgment. SV-FEND (Qi et al., 2023) exploits cross-modal correlations to select the most informative features and uses social context information for detection. FakingRecipe (Bu et al., 2024) detects misinformation by simulating the creation process of audio and video. TikTec (Shang et al., 2021) employs speech text-guided visual object features and MFCC-guided speech textual features. However, these methods are typically designed for general domains and often underperform in domain generalization scenarios.
@@ 3. Proposed Method: DOCTOR
In this section, we propose an innovative DOmain generalization model via ConsisTency and invariance learning for shORt-video misinformation detection (named DOCTOR).
@@ 3.1.2. Overview of DOCTOR
As illustrated in Figure 3, which demonstrates the domain generalization for the short-video misinformation detection, the proposed model, DOCTOR, is designed to address the generalization gap between source and target domains.
@@ 3.2. Cross-modal Interpolation Distillation
If the loss of mixed representation $\hat{X}_{a,v}$ can be optimized to lower values, we can achieve the best state in all modalities at the same time. However, the amount of computation in multi-modal learning is often huge, and it is often difficult to achieve the optimal state, which is impractical in practical applications. Thus, we adopt the cross-modal distillation technique to distill the knowledge in the mixed representation into each modality and then optimize the learned representation to achieve the best state in all modalities. Specifically, for each modal $k\in{(a:Audio,v:Video)}$, we initialize the student model $\gamma_{t}^{k}$ and utilize Simple Moving Average (SMA) to create the teacher model $\hat{\gamma_{t}^{k}}$, which is used to generate more stable and generalizable representations. The $\hat{\gamma}_{t}^{k}$ represents as:
@@
where $\gamma_{t}^{k}$ means that iteration $t$ is the student mode state of modal $k$, and $t_{0}$ is the number of iterations at which to start the SMA. Therefore, the representation produced by the teacher model $\hat{\gamma}_{t}^{k}$ is denoted as $\dot{X}_{k}$, which serves as a stable reference for the modality $k$. We obtain a new cross-modal representation mixup $\dot{X}_{a,v}$ as follows:
@@
where $B$ is the sample set in the training batch, $\tau$ is the temperature coefficient, $\exp(\cdot/\tau)$ is a similarity scaling function used to adjust the discrimination between positive and negative samples, and $\hat{X}_{a}^{i},\hat{X}_{v}^{i}$ are the audio and video embedding vectors of sample $i$.
@@ 3.3. Cross-modal Invariance Fusion
Toward this issue, we propose Cross-modal Invariance Fusion. It designs a novel mechanism for performing core extraction for video features from spatial and temporal perspectives. With extracted multi-modal features, we add noise to them step by step based on the diffusion model to filter out redundant details, thereby underlining the core features of multi-modal. Subsequently, by gradually mitigating the noise influence, we further outline the core information of multi-modal features through cross-modal mutual guidance, enhancing domain invariant features. We aim to mitigate the impact of domain bias and retain domain invariance to enhance the domain generalization ability for misinformation detection.
3.3.1. Multimodal Feature Fusion
@@
(9) $X_{time}=\text{MEAN}(\text{Self-Att}([Seg^{1},Seg^{2},...,Seg^{i}]).$
3.3.2. Diffusion Model Filtering Bias
In the forward diffusion process, we denote the initial distribution of each fusion feature as $X_{emo}^{0}$ (emotion feature), $X_{sem}^{0}$ (semantic feature), $X_{spa}^{0}$ (spatial feature), $X_{time}^{0}$ (temporal feature). For each feature, we gradually incorporate noise sampled from a Gaussian distribution into the true distribution. The steps are as:
@@ 3.3.2. Diffusion Model Filtering Bias
where the equation follows the Markov chain principle and represents the distribution of $X_{F}^{k}$ under $X_{F}^{k-1}$, $\beta_{F}^{k}$ is a weight and indicates the proportion of noise added at the $k$-th step. $I$ represents the identity matrix, which is used to control the covariance structure of the noise. $F$ is a feature of emo feature, sem feature, spa feature, and time feature. $X_{F}^{k}$ can be sampled through the following process:
@@
where $\epsilon$ is the noise sampled from $\mathcal{N}(0,I)$. The overall noise injection process is as follows, filtering out redundant biases:
@@
Table 2. Comparison of general domain misinformation detection performance on FakeSV and FakeTT datasets.
@@
where $\mu_{\theta}$ is the predicted mean (the direction of denoising), and $\Sigma_{\theta}(k)$ is the covariance of the noise. This process gradually generates $X_{F}^{k-1}$, $X_{F}^{k-2}$, … , $X_{F}^{0}$ from $X_{F}^{k}$ that is, gradual denoising.
4. Experiments
4.1. Experimental Setups
4.1.1. Datasets
4.1.2. Baselines
We selected some popular misinformation short-video detection methods, such as FakingRecipe (Bu et al., 2024), OpEvFake (Zong et al., 2024), TikTec (Shang et al., 2021), HCFC-Hou (Hou et al., 2019), HCFC-Medina (Serrano et al., 2020), FANVM (Choi and Ko, 2021), SV-FEND (Qi et al., 2023). We also compared with the novel general domain generalization method CMRF (Fan et al., 2024) and the text-image misinformation domain generalization methods MMDFND (Tong et al., 2024) and MDFEND (Nan et al., 2021). We also select large language models, such as GPT-4 (Achiam et al., 2023) and GPT-4V (Yang et al., 2023b) for comparison.
4.2. Experimental Results
4.2.1. Domain generalization experiments
@@ 4.2.3. Ablation experiments
Table 4. Learning rate for DOCTOR model (Accuracy).
Dataset 1e-5 5e-5 1e-4 5e-4 1e-3 5e-3
FakeSV 82.32 84.91 83.45 80.78 79.53 78.44
FakeTT 72.24 74.31 76.86 78.57 78.92 77.39
4.2.4. Parameter sensitivity analysis
@@ 5. Conclusion
In this paper, we firstly explore the shortcomings of the domain generalization of short-video misinformation detection, which are mainly divided into two aspects: (1) Varying Modal-dependence in Different Domains and (2) Domain-bias Enhancement in Cross-modal Fusion. To bridge these gaps, we propose an innovative DOmain generalization model via ConsisTency and invariance learning for shORt-video misinformation detection (named DOCTOR). It incorporates two core modules: (a) Cross-modal Interpolation Distillation and (b) Cross-modal Invariance Fusion. We validate the domain generalization effectiveness of misinformation detection of DOCTOR through comprehensive experiments. In addition, in-depth analysis, including ablation studies, demonstrates the design rationality of each module in the DOCTOR model.
Acknowledgements.